the concession stand at an ice hockey rink had receipts of $7200 from selling a total of 3000 sodas and hot dogs. if each soda sold for $2 and each hot dog sold for $4, how many of each were sold?
Answer by Cromlix(4381)

Sodas = x
Hot dogs = y
x + y = 3000.....1
2x + 4y = 7200.....2
Multiply Eq (1) by 4
x + y = 3000.....1
4x + 4y = 12000.....3
2x + 4y = 7200......2
Subtract Eq(2) from Eq(3)
2x = 4800
x = 2400
Substitute x = 2400 into
Eq(1)
x + y = 3000.......1
2400 + y = 3000
y = 3000 - 2400
y = 600
Therefore 2,400 sodas and
600 hot dogs were sold.
Hope this helps.
:-)